# Webly-Supervised Image Manipulation Localization via Category-Aware Auto-Annotation

### v1

@@ VII-B Experiments on the Web-Supervision Method (IML Task)
+MIMLv1 (Conf.) .755 0% .794 0% .385 6% .458 10% .352 9% .408 8% .596 58% .678 57% .340 13% .401 14% .486 18% .548 19%
+MIMLv2 .767 2% .794 0% .387 6% .458 10% .407 21% .472 20% .618 60% .696 58% .370 20% .422 18% .510 22% .568 22%
+MIMLv2+O.J .770 2% .811 2% .466 22% .511 20% .439 27% .502 25% .619 60% .700 59% .309 4% .357 3% .521 24% .576 23%
@@
TABLE VIII: Comparison study of CIML on SPG (left) and SDG (right). ’*’ denotes model trained with both SDG and SPG data. ’Nonzero’ denotes using the non-zero region of the difference map as mask, ’OTSU’ denotes the difference map binarized by OTSU. ’QES’ denotes the predictions filtered by our QES metric.
@@
TABLE IX: Ablation study of CIML on SPG. ’Forged’ denotes using the forged image as input, ’Auth.’ denotes using the authentic image as input, ’Diff.’ denotes using the absolute difference map between the forged and authentic image as input.
@@ VII-C Experiments on the CAAAv2 Models (CIML task)
This subsection evaluates the effectiveness of the proposed CAAAv2 in the CIML task by comparing its DASS and Corr-DINO models against other CIML models on SPG and SDG test datasets, respectively. We also conduct detailed ablation studies for in-depth analysis.
@@
Implementation Details. Given that real-world forgeries in the IMD20 [32] dataset closely resemble the target images for annotation, we use a portion of this dataset to evaluate model performance using IoU and binary F1-score metrics. We categorize the forged images in IMD20 into SPG or SDG and randomly split them into training and test sets with a 3:1 ratio. CASIAv2 [31] and approximately 1M images synthesized from the COCO [49] dataset are also used for training. Input images are resized to $\times512$, and a consistent training pipeline is applied across all methods.
@@
Comparison Study. Table VIII presents comparison results for SPG (left) and SDG (right) separately. Our method clearly outperforms these previous methods, primarily because our category-aware paradigm significantly reduces task difficulty. As shown in the right side of Table VIII, our Corr-DINO, prior to QES filtering, achieves an IoU of 0.744. This is 4.2 points higher than the SACM [14] (our conference version), demonstrating that our frozen-denoising paradigm notably alleviates the overfitting issue when training on limited data. Notably, DMVN [26] and DMAC [51] models trained on both SPG and SDG data perform worse than those trained exclusively on one category. This is attributed to the confusion caused by similar backgrounds in SPG image pairs.
@@
Ablation Study for the DASS Model Input. Table IX presents the results, confirming that both the image pair and the image difference map are essential for accurate prediction in SPG. The image difference map between forged and authentic images can roughly indicate the forged regions, while the images themselves provide semantic information to help the model denoise the difference map.
@@
Ablation Study for the Corr-DINO Model Backbone. The results are shown in Table X, settings (1-19). For the ConvNeXt CNN model, the trainable version performs better. For all ViT models, however, frozen versions perform significantly better. Since ViT models tend to overfit with limited training data, freezing the backbone alleviates catastrophic forgetting and overfitting. LoRA [53] fine-tuned versions exhibit similar performance to their frozen counterparts, as LoRA also mitigates forgetting. The frozen DINOv2-Base [13] backbone performs optimally and is therefore adopted.
@@
Ablation Study of the Corr-DINO Model Modules. The results are shown in the bottom of Table X, setting (20) is the baseline without any proposed modules, setting (24) is the full model with all proposed modules. In settings (21, 22, 23), we remove the proposed Multi-Aspect Denoiser, Feature Super Resolution and Learnable Aggregation modules from the full model (24) respectively. Setting (24) significantly outperforms (21). This validates that correlation features extracted from the frozen DINOv2 are highly noisy, and that denoising through multi-aspect feature integration is essential for accurate predictions. Setting (24) also significantly outperforms (22). This confirms that initial correlation features are too coarse for precise mask predictions, and our Feature Super Resolution module effectively addresses this by reconstructing fine-grained details. Finally, Setting (24) is superior to (23) because our learnable aggregation mitigates information loss inherent in rule-based aggregation.
@@
Ablation Study of the QES. The last two rows of Table X show the results. The results of setting (25) are obtained by evaluating only the predictions from setting (24) where QES>>0.5. QES-filtered predictions from setting (25) achieve an IoU of 0.912, representing a 16.8 point increase over setting (24). This confirms that QES effectively filters out unsatisfactory predictions, thereby notably improving annotation quality. Table XI confirms QES’s high relevance to the IoU metric without relying on ground truth for evaluation.
@@
Our Models are Adequate Auto-Annotators. In Table VIII, both our DASS and Corr-DINO achieve high IoU scores. Considering the errors in IMD20’s ground truth [54], especially in SPG images, our actual accuracy is higher. Clearly, our methods are sufficient for obtaining accurate auto-annotations.
@@
TABLE X: Ablation study of CIML on SDG. ‘Freeze‘ denotes freezing the backbone and ‘L‘ denotes LoRA [53] fine-tuning. ‘MAD‘ denotes the Multi-Aspect Denoiser. ‘FSR‘ denotes the Feature Super Resolution. ‘LA‘ denotes the Learnable Aggregation. ‘QES‘ denotes filtering the prediction with our QES.
@@
TABLE XII: Image manipulation localization performance on SACP [58] dataset. ‘Forgery Baseline‘ denotes using the tampCOCO [7] and CASIAv2 [31] datasets for model pretraining
@@
(1) DFCN [59] Image-Net [60] .466 .607
(2) MVSS-Net [20] Image-Net [60] .401 .534
(3) PSCC-Net [37] Image-Net [60] .482 .620
(4) RRU-Net [43] Image-Net [60] .517 .651
(5) CFL-Net [61] Image-Net [60] .433 .571
(6) DTD [62] Image-Net [60] .588 .712
(7) TIFDM [63] Image-Net [60] .576 .703
(8) Web-IML (Ours) Image-Net [60] .773 .862
@@
(12) Web-IML (Ours) (10) + PSCC Synthetic [37] .762 .853
@@ VII-D Experiments on the Downstream Document IML Task
We fine-tune our Web-IML under various pretraining settings on the handcrafted document IML benchmark SACP [58], using the same training and test configurations as prior works [63]. Our Web-IML notably outperforms prior methods. Moreover, Table XII shows that settings (11, 12, 13), which incorporate additional training data from DEFACTO [8], PSCC [37], and OpenSDID [48] into the baseline (Setting 10), perform worse than the baseline. This confirms that simply increasing the scale and diversity of pretraining data does not lead to better generalization. In contrast, the model pretrained on our web supervision data (15) shows a significant 2.2 point IoU improvement over the baseline (10), and is 0.8 points higher than the model pretrained on MIMLv1 (14) from our conference paper. These results demonstrate that the cheap data from our web supervision method is of sufficient quality to enable models to learn semantic-agnostic forensic features and generalize across unseen real-world IML tasks.
VIII Conclusion

### v2

@@ VI-B Self-Rectification Module
+WSCL [66] .758 1% .797 1% .362 0% .410 0% .326 1% .373 1% .307 24% .352 22% .310 5% .359 4% .412 4% .458 3%
+MIMLv1 (Conf.) .755 0% .794 0% .385 6% .458 10% .352 9% .408 8% .596 158% .678 157% .340 13% .401 14% .486 18% .548 19%
+MIMLv2 .767 2% .794 0% .387 6% .458 10% .407 21% .472 20% .618 160% .696 158% .370 20% .422 18% .510 22% .568 22%
+MIMLv2+O.J .770 2% .811 2% .466 22% .511 20% .439 27% .502 25% .619 160% .700 159% .309 4% .357 3% .521 24% .576 23%
SparseViT B.L. .702 0 .747 0 .320 0 .352 0 .288 0 .343 0 .199 0 .239 0 .177 0 .232 0 .337 0 .383 0
+PSCC-Synth. .627 11% .677 9% .304 5% .333 5% .256 11% .311 9% .196 1% .243 1% .169 5% .232 0% .310 8% .359 6%
+DEFACTO[20] .692 1% .730 2% .257 20% .283 20% .310 8% .364 6% .186 7% .225 6% .155 12% .198 15% .320 5% .362 6%
+AIGC-Synth. .533 24% .583 22% .119 63% .141 60% .133 54% .174 49% .088 56% .110 54% .033 81% .045 81% .181 46% .211 45%
+WSCL [66] .709 1% .755 1% .324 1% .356 1% .301 4% .352 3% .267 34% .314 31% .184 4% .241 4% .357 6% .404 6%
+MIMLv1 (Conf.) .714 2% .762 2% .394 23% .458 30% .306 6% .359 5% .535 169% .617 158% .215 21% .270 16% .433 28% .493 29%
+MIMLv2 .723 3% .771 3% .441 38% .509 45% .338 17% .396 15% .544 173% .624 161% .234 32% .289 25% .456 35% .518 35%
+MIMLv2+O.J .720 3% .767 3% .477 49% .547 55% .346 20% .409 19% .551 178% .633 165% .252 42% .325 40% .469 40% .536 40%
IMDPrompter B.L. .718 0 .765 0 .463 0 .522 0 .346 0 .419 0 .261 0 .310 0 .354 0 .414 0 .428 0 .486 0
+PSCC-Synth. .682 5% .729 5% .473 2% .531 2% .295 15% .356 15% .262 0% .312 1% .286 19% .337 19% .400 7% .453 7%
+DEFACTO[20] .701 2% .749 2% .439 5% .508 3% .284 18% .347 17% .243 7% .286 8% .281 21% .329 21% .390 9% .444 9%
+AIGC-Synth. .630 12% .678 11% .254 45% .289 45% .218 37% .263 37% .222 15% .264 15% .263 26% .309 25% .317 26% .361 26%
+WSCL [66] .701 2% .750 2% .485 5% .543 4% .360 4% .433 3% .285 9% .336 8% .356 1% .417 1% .437 2% .496 2%
+MIMLv1 (Conf.) .724 1% .769 1% .542 17% .610 17% .418 21% .501 20% .601 130% .696 125% .392 11% .456 10% .535 25% .606 25%
+MIMLv2 .729 2% .773 1% .564 22% .635 22% .426 23% .512 22% .614 135% .708 128% .418 18% .490 18% .550 29% .624 28%
+MIMLv2+O.J .735 2% .778 2% .626 35% .706 35% .433 25% .521 24% .635 143% .721 133% .425 20% .498 20% .571 33% .645 33%
VII Experiments
VII-A Experiments on the Web-IML Model (IML Task)
Implementation Details. For Web-IML, we adopt ConvNeXt-Base [68] as the feature extractor. The model is trained for 200k iterations with a batch size of 16, using a $\times512$ input size, consistent with previous works [18, 49]. We employ cross entropy loss and AdamW [69] optimizer, with the learning rate linearly decaying from 1e-4 to 1e-6. CASIAv2 [46] and tampCOCO [19] serve as baseline training sets. For fair comparison, mask predictions are binarized with a fixed threshold of 0.5 and evaluated using IoU and binary F1-score.
Comparison Study. Table II shows that our Web-IML significantly outperforms existing methods. For example, it achieves an average IoU that is 21.6 points higher than SparseViT. Furthermore, Web-IML consistently outperforms our conference version, APSC-Net, across various settings. These results demonstrate the strong generalization capabilities of our Web-IML model. Qualitative comparison is illustrated in Fig. 10.
Ablation Study. The ablation results are shown in Table III. Setting (2), which includes the proposed Multi-Scale Perception (MP) module, achieves 5.2 points higher IoU than the baseline (1). This indicates that the proposed MP module effectively integrates high-level and low-level features to detect diverse image artifacts. Setting (3), which further incorporates the Self-Rectification (SR) module, yields an additional 3.9 points IoU improvement over Setting (2), as SR enables the model to detect and correct its errors. Setting (4) outperforms (3) via our nested attention module that enables in-depth analysis. Setting (5) outperforms (4) by including a second round of self-rectification refinement, demonstrating that further refinement can minimize incorrect predictions.
VII-B Experiments on the Web-Supervision Method (IML Task)
This subsection evaluates the effectiveness of our proposed web-supervision approach in addressing the data scarcity issue for IML. Our web-supervision encompasses both the MIMLv2 dataset and the Object Jitter method. We compare the performance of the same IML model trained on data created by our web-supervision approach against other synthetic methods and WSCL [66], a representative weakly supervised framework.
Implementation Details. The implementation details remain the same as in the previous subsection VII.A. When supplementing the baseline training sets with additional IML training sets, the total training volume is kept constant with a uniform sampling ratio across all datasets. For models integrated with WSCL, training involved two types of data: direct training with the fully (pixel-level) annotated tampCOCO [19] and CASIAv2 [46] datasets, and WSCL-specific training with weakly (image-level) annotated MIMLv2 and SA [57] dataset images. The training volume and other configurations remain the same as ours.
Comparison Study for the MIMLv2 Dataset. Table V presents the IML performance of the Web-IML model when trained on different datasets. The baseline (1) uses the two most common training sets: tampCOCO [19] and CASIAv2 [46]. Incorporating synthetic training data from PSCC-Net [56], DEFACTO [20] or OpenSDID [21] diminishes the average model performance (settings 2, 3, 4). This is because the first two synthetic datasets do not notably surpass the baseline data in quality or diversity, while the AIGC forgeries from OpenSDID [21] differ significantly from non-AIGC forgeries. These results confirm that simply increasing the scale of training data does not lead to improved performance. In contrast, including the MIMLv1 (setting 6) or MIMLv2 (setting 7) datasets significantly improves model performance by an average of 10 and 12 points, respectively, without increasing training or inference burden. Additionally, although WSCL (setting 5) offers improvements over baselines, its effectiveness is considerably lower than that of our web-supervision. This is because our method explicitly produces accurate mask annotations, by leveraging the image correspondence that substantially reduces task difficulty. Conversely, WSCL lacks the ability to leverage this valuable information, and relies heavily on unstable noise domain features. These results confirm that our MIMLv1 and MIMLv2 datasets considerably alleviate data scarcity and overfitting issues in real-world scenarios. This is attributable to their large scale, manual forgery, and considerably higher quality and diversity compared to other datasets. Notably, MIMLv2 yields greater improvement due to its larger quantity and superior diversity compared to MIMLv1.
Comparison Study for the Object Jitter. Table V shows that including the training data generated by our Object Jitter method further improves the model when combined with MIMLv1 (setting 8 outperforms 6) or MIMLv2 (setting 10 outperforms 7) datasets. This improvement is attributed to Object Jitter’s consistent generation of high-quality hard samples, which effectively supplements the training data by compensating for the relative scarcity of copy-move forgeries.
Ablation Study for the Object Jitter. Table VI presents the ablation study for Object Jitter operations and image source. Setting (1) is the baseline, where Web-IML is trained with MIMLv2 but without Object Jitter. Setting (7) represents the full Object Jitter method. Settings (2, 3, 4) involve removing the size jitter, texture jitter, and exposure jitter operations, respectively; all of them result in worse performance than setting (7). These results demonstrate that all three operations contribute to improved performance through increased data diversity. Notably, Setting (2), which lacks size jitter, is significantly worse than (7). Meanwhile, Setting (5), which uses only size jitter, performs close to (7). This indicates that the size jitter operation is the primary contributor to the improvement. Since most Internet forgeries are not generated by copy-move, they exhibit exposure and texture artifacts. The size jitter operation, without altering the object’s exposure or texture, reduces the model’s reliance on exposure or texture cues, thus better supplementing MIMLv2. Furthermore, Setting (6) applies the full Object Jitter method using COCO [67] as the image source, and setting (7) uses 400k random images from SA [57]. The similar performance of Settings (6) and (7) demonstrates that the Object Jitter method is robust to the choice of image source.
Verification of Conclusions on Other IML Models. We extended our verification process to other IML models. Specifically, we evaluated MIMLv2 and Object Jitter on PSCC-Net [56], CAT-Net [19], TruFor [49], SparseViT [18] and IMDPrompter [70] (a large model based on SAM [57]) using the same training data and configuration as ours. The results in Table VII clearly indicate that MIMLv1, MIMLv2, and Object Jitter all lead to significantly higher scores. MIMLv2, in particular, consistently yields significantly greater improvement than MIMLv1 (conference version). Moreover, Object Jitter further boosts performance when integrated with MIMLv2. These results robustly confirm our previous analyses and conclusions.
@@ VII-B Experiments on the Web-Supervision Method (IML Task)
TABLE VIII: Comparison study of CIML on SPG (left) and SDG (right). ’*’ denotes model trained with both SDG and SPG data. ’Nonzero’ denotes using the non-zero region of the difference map as mask. ’QES’ denotes the predictions filtered by our QES metric.
@@
TABLE IX: Ablation study of CIML on SPG. ’Forged’ denotes using the forged image as input, ’Auth.’ denotes using the authentic image as input, ’Diff.’ denotes using the difference map between the forged and authentic image as input.
@@ VII-C Experiments on the CAAAv2 Models (CIML task)
This subsection evaluates the effectiveness of the proposed CAAAv2 in the CIML task by comparing its DASS and Corr-DINO models against other CIML models. A higher CIML score means more accurate auto-annotation. We also conduct detailed ablation studies for in-depth analysis.
@@
Implementation Details. Given that real-world forgeries in the IMD20 [45] dataset closely resemble the target images for annotation, we use a portion of this dataset to evaluate model performance using IoU and binary F1-score metrics. We categorize the forged images in IMD20 into SPG or SDG and randomly split them into training and test sets with a 3:1 ratio. CASIAv2 [46] and synthetic COCO [67] datasets are also used for training. Input images are resized to $\times512$, and a consistent training pipeline is applied across all methods.
@@
Comparison Study. Table VIII presents comparison results for SPG (left) and SDG (right) separately. Our method clearly outperforms these previous methods because our category-aware and prior-feature-denoising paradigm significantly reduces task difficulty. As shown in the right side of Table VIII, our Corr-DINO, prior to QES filtering, achieves an IoU of 0.744. This is 4.2 points higher than the SACM [24] (our conference version), demonstrating that our frozen-denoising paradigm notably alleviates the overfitting issue when training on limited data. Notably, DMVN [40] and DMAC [71] models trained on both SPG and SDG data perform worse than those trained exclusively on one category. This is attributed to the confusion caused by similar backgrounds in SPGs, confirming the necessity of our category-aware approach.
@@
Ablation Study for the DASS Model Input. Table IX presents the results, confirming that both the prior feature difference map and the image pair are essential for accurate prediction in SPG. The difference map between forged and authentic images can roughly indicate the forged regions, while the images themselves provide semantic information to help the model denoise the difference map.
@@
Ablation Study for the Corr-DINO Model Prior Feature Extractor. The results are shown in Table X, settings (1-24). For ConvNeXt, the trainable version performs better. For all ViT-based models, however, frozen versions perform significantly better. Since ViT models tend to overfit with limited training data, freezing the backbone alleviates catastrophic forgetting and overfitting. LoRA [73] fine-tuned versions exhibit similar performance to their frozen counterparts, as LoRA also mitigates forgetting. The frozen DINOv2-Base [42] backbone performs optimally and is therefore adopted.
@@
Ablation Study of the Corr-DINO Model Modules. The results are shown in the bottom of Table X, setting (25) is the baseline without any proposed modules, setting (29) is the full model with all proposed modules. In settings (26, 27, 28), we remove the proposed Multi-Aspect Denoiser, Feature Super Resolution and Learnable Aggregation modules from the full model (29) respectively. Setting (29) significantly outperforms (26). This validates that correlation features extracted from the frozen DINOv2 are highly noisy, and that denoising through multi-aspect feature integration is essential for accurate predictions. Setting (29) also significantly outperforms (27). This confirms that initial correlation features are too coarse for precise mask predictions, and our Feature Super Resolution module effectively addresses this by reconstructing fine-grained details. Finally, Setting (29) is superior to (28) because our learnable aggregation mitigates information loss inherent in previous rule-based aggregation.
@@
Ablation Study of the QES. The last two rows of Table X show the results. The results of setting (30) are obtained by evaluating only the predictions from setting (29) where QES>>0.5. QES-filtered predictions from setting (30) achieve an IoU of 0.912, representing a 16.8 point increase over setting (29). This confirms that QES effectively filters out unsatisfactory predictions, thereby notably improving annotation quality. Table XI confirms QES’s high relevance to the IoU metric without relying on ground-truth for evaluation.
SPG/SDG Classification. We evaluate the performance of the SPG/SDG classification model within the proposed CAAAv2 on IMD20 [45] dataset. Considering that a very small proportion of SPG image pairs are not spatially aligned, which could negatively impact the prediction’s quality, we also include an additional linear binary classification layer to filter them out. The classification results are presented in Table XII. It is evident that the voting ensemble of the classification models we used is accurate enough for SPG/SDG classification, as analyzed in Section III.A. In practice, to further ensure accuracy, we only keep images with a classification confidence above 99% and discard the rest. Our QES also effectively filters the annotations of potentially misclassified samples.
@@
Our Models are Adequate Auto-Annotators. In Table VIII, both our DASS and Corr-DINO achieve high IoU scores. Considering the errors in IMD20’s ground-truth [74], especially in SPG images, our actual accuracy is higher. Clearly, our methods are sufficient for obtaining accurate auto-annotations.
@@
TABLE X: Ablation study of CIML on SDG. ‘Freeze‘ denotes freezing the backbone and ‘L‘ denotes LoRA [73] fine-tuning. ‘MAD‘ denotes the Multi-Aspect Denoiser. ‘FSR‘ denotes the Feature Super Resolution. ‘LA‘ denotes the Learnable Aggregation. ‘QES‘ denotes filtering the prediction with QES.
@@
(20) DINOv3-B [78] $\times$ $\times$ .692 .588 .636 .478
(21) DINOv3-B [78] L $\times$ .690 .753 .720 .598
(22) DINOv3-B [78] $\times$ .709 .787 .746 .605
(23) DINOv3-L [78] $\times$ $\times$ .718 .593 .650 .496
(24) DINOv3-L [78] $\times$ .722 .791 .755 .621
@@
TABLE XII: Classification results for SDG, SPG and misaligned SPG. ‘Ensemble‘ denotes the ensemble of the models.
method SDG SPG Misaligned
P R F P R F P R F
DiNAT [79] .992 .992 .992 1 .996 .998 .867 .929 .897
SwinTrans [80] .996 .981 .988 1 .996 .998 .737 1 .849
ConvNeXt [68] .996 .992 .994 1 .996 .998 .875 1 .933
Ensemble .996 1 .998 1 .996 .998 1 1 1
TABLE XIII: Image manipulation localization performance on SACP [81] dataset. ‘Forgery Baseline‘ denotes using the tampCOCO [19] and CASIAv2 [46] datasets for model pretraining
@@
(1) DFCN [82] ImageNet [83] .466 .607
(2) MVSS-Net [35] ImageNet [83] .401 .534
(3) PSCC-Net [56] ImageNet [83] .482 .620
(4) RRU-Net [60] ImageNet [83] .517 .651
(5) CFL-Net [84] ImageNet [83] .433 .571
(6) DTD [85] ImageNet [83] .588 .712
(7) TIFDM [86] ImageNet [83] .576 .703
(8) Web-IML (Ours) ImageNet [83] .773 .862
@@
(12) Web-IML (Ours) (10) + PSCC-Synthetic [56] .762 .853
@@ VII-D Experiments on the Downstream Document IML Task
We fine-tuned the Web-IML model, pretrained under various configurations, on the handcrafted real-world document IML benchmarks, SACP [81] and RTM [88], adopting the same training and test configurations as previous studies [86, 88]. Across both benchmarks, our Web-IML model consistently and notably outperforms existing methods. A key observation from Tables XIII and XIV is that settings 11, 12, and 13, which augmented the baseline (Setting 10) with additional pretraining data from DEFACTO [20], PSCC [56], and OpenSDID [21], respectively, exhibited degraded performance. This confirms that merely increasing the scale and diversity of pretraining data does not lead to improved real-world generalization. In contrast, the model pretrained on our web-supervision data (Setting 15) demonstrates significant improvement. It achieved 2.2 points higher IoU on SACP and 4.7 points on RTM compared to the baseline (Setting 10). Furthermore, it also surpassed the model pretrained on MIMLv1 (Setting 14) by 0.8 points higher IoU on SACP and 1.5 points on RTM. These results demonstrate that the cost-effective data created by our web-supervision method possesses sufficient quality and relevance to enable models to learn robust semantic-agnostic forensic features, thereby generalizing effectively across diverse and unseen real-world IML tasks.
TABLE XIV: Image manipulation localization performance on RTM [88] dataset. ‘Forgery Baseline‘ denotes using the tampCOCO [19] and CASIAv2 [46] datasets for model pretraining.
Num. Methods Pretraining data IoU F1
(1) RRU-Net [60] ImageNet [83] .037 .072
(2) PSCC-Net [56] ImageNet [83] .033 .064
(3) MVSS-Net [35] ImageNet [83] .051 .097
(4) CAT-Net [19] ImageNet [83] .113 .203
(5) DTD [85] ImageNet [83] .065 .122
(6) Mask2Former [89] ADE20k [87] .124 .172
(7) ASCFormer [88] ADE20k [87] .198 .329
(8) Web-IML (Ours) ImageNet [83] .192 .323
(9) Web-IML (Ours) ADE20k [87] .207 .343
(10) Web-IML (Ours) Forgery Baseline .194 .325
(11) Web-IML (Ours) (10) + DEFACTO [20] .186 .313
(12) Web-IML (Ours) (10) + PSCC-Synthetic [56] .189 .318
(13) Web-IML (Ours) (10) + AIGC Synthetic [21] .185 .311
(14) Web-IML (Ours) (10) + MIMLv1 [24] .226 .369
(15) Web-IML (Ours) (10) + Web Supervision (Ours) .241 .389
Figure 10: Qualitative evaluation on image manipulation localization. From left to right: input image, Ground-Truth (abbreviated as GT), RRU-Net (RRU), PSCC-Net (PSCC), MVSS-Net (MVSS), IF-OSN (OSN), CAT-Net (CAT), TruFor (T.F.), SparseViT (S.ViT), APSC-Net (APSC) and ours.
Figure 11: Both Web-IML and previous methods sometimes struggle to produce precise mask when the forged region exhibits few visual artifacts. From left to right: input image, Ground-Truth (abbreviated as GT), RRU-Net (RRU), MVSS-Net (MVSS), IF-OSN (OSN), CAT-Net (CAT), TruFor (T.F.), SparseViT (S.ViT), APSC-Net (APSC) and ours.
VIII Limitations
Although our QES metric is highly effective in ensuring overall annotation quality, a small number of correct annotations may also be filtered out. In addition, our Web-IML is designed to maximize accuracy by leveraging web-scale supervision with a medium-sized model (140M). As a result, it is not highly computationally efficient (achieving 10 frames per second on a single 3090 GPU). Figure 11 presents some failure cases, showing that both Web-IML and previous methods may struggle to produce precise masks when the manipulated region exhibits very few visual artifacts.
IX Conclusion